# Optical Physiological Signals

### v1

@@ 6 Conclusion and Future Work
Our results show that PaPaGei outperforms baselines in 14 tasks, with classification and regression improvements of 4.7%-6.3% and 2.9%-4.9%, respectively. PaPaGei-S excelled in cardiovascular tasks like BP, Hypertension, and HR, which can be attributed to the sVRI and IPA objectives, and PaPaGei-P outperformed baselines like Moment, excelling in tasks such as Smoking and Arousal. Ablation studies confirmed that the model with all three SSL objectives performs best, with sVRI highlighted as a key component and IPA and SQI providing positive knowledge transfer in multi-task setups. PaPaGei is both data- and size-efficient (5M), making it ideal for medical applications where large models (200M+) are impractical due to on-device limitations or data privacy concerns with cloud model inference. While combining PaPaGei-P and PaPaGei-S objectives into one model might seem intuitive, it is impractical because it would constrain positive pairs on both sVRI and the number of participants, resulting in too many unique labels with limited samples per label. Our case studies also showed that PaPaGei-S captured personal medical information due to well-dispersed embeddings, compared to baselines. Future work should focus on diversifying training data, investigating sampling rate effects, and exploring multi-modal approaches or alternative architectures. In conclusion, PaPaGei represents a significant advancement in foundation models for analyzing PPG signals in resource-constrained medical environments, with its open-source nature encouraging further research and development in healthcare applications.
@@
Models, data, and code will be made publicly available upon release, ensuring accessibility for future research. Our study exclusively utilizes publicly accessible datasets, which can be requested or downloaded from the respective study group websites, allowing others to easily obtain the data for their own analyses. In Section 4 and Appendix §B, we provide comprehensive descriptions of the datasets, ground-truth annotations, and data pre-processing methods used in our experiments, ensuring transparency in our data handling procedures. The code to run our model will be available with user-friendly examples. We have provided a detailed overview of the model architecture and its hyperparameters in Sections 3, 4.1, and Appendix §A. With these detailed resources, our work is designed to be reproducible, enabling future research to build upon our findings.
@@ Appendix A Training and Inference Details
Architecture & Pre-training. The architecture of our ResNet 18-block encoder is described in Tables 7, 7, and 7. Each 1D convolution layer is configured with a kernel size of 3 and a stride of 2, while the max-pooling layer utilizes a kernel size of 3 with a stride of 1. We start with a filter size of 32, which doubles every 4 blocks to capture progressively more complex features. Dropout is applied with a probability of 0.5 to prevent overfitting. This backbone architecture is used across different methods in our experiments to ensure consistency and make for a fair comparison during evaluation. In PaPaGei-S, we project the learned representations to a 512-dimensional embedding after the convolutional block (we investigated larger embedding sizes of 768 and 1024 and found no significant performance changes). This embedding is then passed through two Mixture of Experts (MoE) blocks, each containing three experts. Each expert block consists of two sequential linear layers, with sizes 256 and 1, which are used for IPA and SQI prediction tasks. It is noteworthy that both BYOL and TF-C require multiple encoders and different projection heads, resulting in variations in model sizes. For these methods, we use existing implementations available online [^3] https://github.com/chengding0713/SiamQuality [^4] https://github.com/mims-harvard/TFC-pretraining, but apply our encoder as the backbone to ensure consistency. Our models are pre-trained for 15,000 steps using the Adam optimizer, with a learning rate of $10^{-4}$. We use a batch size of 128 for training since after various trials we did not observe significant differences in performance with batch sizes of 64 and 256. Additionally, we did not perform an exhaustive evaluation of different augmentation settings but instead used transformations and values based on prior research (Abbaspourazad et al., 2023; Tang et al., 2020). For model training, we primarily used PyTorch (Paszke et al., 2019). The NTXentLoss implementation was sourced from the PyTorch Metric Learning package [^5] https://github.com/KevinMusgrave/pytorch-metric-learning.
@@ Appendix B Datasets and Tasks
Table 8: The task evaluation benchmark of PaPaGei. Datasets highlighted in grey are unseen during training, thus, the corresponding tasks are out-of-domain. The rest were used for pre-training but their test sets and labels are held out. For task type, B/M/R refer to Binary classification, Multi-class classification (#classes), and Regression, respectively.
#ID Dataset SR (Hz) Collected by Task Task Type #Subjects (#Samples)
@@
VitalDB. The VitalDB dataset provides comprehensive monitoring of vital signs and physiological parameters from 6,388 surgical cases. This high-resolution dataset includes a wide range of intraoperative monitoring variables such as heart rate, blood pressure, oxygen saturation, and other critical physiological signals, collected at frequent intervals throughout surgery. The surgical operation belongs to one of the eleven categories: colorectal, biliary/pancreas, stomach, major resection, minor resection, breast, transplantation, thyroid, hepatic, vascular, and others. After the data cleaning process, we narrowed the dataset down to 5,866 subjects with complete and usable information. As depicted in Figure 10, we observe that the gender distribution is relatively balanced, with nearly equal representation of male and female patients. Additionally, the majority of the subjects fall within the age range of 50 to 70, with a significant proportion being around 60 years old. The ICU label corresponds to whether the person was admitted to the ICU or not.
@@
MIMIC-III. In our analysis, we utilize the MIMIC-III waveform database matched subset, which comprises data from 10,282 ICU patients. From this dataset, we focus specifically on extracting photoplethysmogram (PPG) data, provided it is available for each patient. To ensure the quality of the data, we set a criterion of at least 1 minute of usable PPG signal that must be present. After performing a thorough data cleaning process, we end up with a cohort of 5,596 subjects with reliable PPG data. As illustrated in Figure 11, the dataset shows a gender imbalance, with a higher proportion of male patients compared to female patients. Additionally, the majority of subjects are aged 60 years or older, reflecting a typical ICU population that often includes elderly patients with critical health conditions.
@@
MESA. The Multi-Ethnic Study of Atherosclerosis (MESA) sleep sub-study gathered data from 2,237 participants through overnight, unattended polysomnography to assess various sleep parameters. After the data cleaning process, we retained 2,055 subjects for analysis. As shown in Figure 12, the dataset shows a slightly larger proportion of female participants. The age distribution reveals that most subjects are between 60 and 80 years old, reflecting an older adult population, which is commonly studied concerning sleep disorders and cardiovascular risks.
In this study, we use the Apnea-Hypopnea Index (AHI) with at least 3% and 4% oxygen desaturation as the primary measure for diagnosing sleep apnea, as recommended by the American Academy of Sleep Medicine (AASM) (Ruehland et al., 2009). These thresholds indicate the severity of sleep apnea, with oxygen desaturation during apneas/hypopneas being a critical factor. We predict these AHI values directly in our regression models. Additionally, we classify participants with any history of smoking as smokers. This approach allows us to account for both current and former smokers, capturing a broader range of smoking-related health risks within our analysis.
@@
VitalVideos (VV) (Skin Tone). The Vital Videos study is an ongoing project that collects data on vital signs, videos, and blood pressure across a variety of conditions, including variations in lighting, background, and skin tone. For our analysis, we used data from two groups, totalling 231 subjects, from Europe and Sub-Saharan Africa. As shown in Figure 14, most subjects have a Fitzpatrick skin tone of 5 or 6, indicating darker skin. The dataset is primarily composed of female participants, with an age range between 40 and 60 years. Additionally, the majority of subjects had a systolic blood pressure of around 125 and a diastolic pressure of around 80, suggesting that most individuals in the study were relatively healthy.
@@
PPG-BP. The PPG-BP consists of short PPG recordings from 219 subjects collected at 1000Hz. For each subject, there are three 2.1s recordings. For our analysis, we zero pad them to 10s. In Figure 15, the age distribution shows that most subjects are between 40 and 80 years old, with fewer participants under 40. Furthermore, the majority of individuals have hypertension. In terms of gender, the dataset has slightly more females than males. The distribution of systolic blood pressure is centered around 120-140, indicating a population with normal to moderately elevated blood pressure, while diastolic blood pressure predominantly falls between 70 and 90. Lastly, the average heart rate for most subjects ranges between 70 and 90 beats per minute.
@@
ECSMP. The ECSMP dataset was gathered to study the relationship between emotion, cognition, and sleep in 89 subjects. As shown in Figure 17, the majority of the participants are young adult females, with an average age of around 25 years. Mood disturbances were measured using the Profile of Mood States (POMS) scale, which captures various aspects of emotional states. To classify participants into high versus low mood disturbance categories, we binarized the Total Mood Disturbance (TMD) values by using the median as the cutoff point.
@@
WESAD. The wearable stress and affect detection dataset is a multi-modal dataset collected from 15 subjects using various sensor modalities. In this study, participants were exposed to videos designed to elicit different affective states, such as amusement, meditation, stress, and baseline conditions. Following each session, participants completed the Self-Assessment Manikins (SAM) questionnaire (Bradley & Lang, 1994), which provided the ground-truth values for valence and arousal. In our analysis, we binarized these values by categorizing valence and arousal as low (1) when less than 5 and high (0) otherwise. Then, we perform regression at the segment level. As shown in Figure 18, arousal levels are generally low, while valence tends to be high in most cases.
@@
PPG-DaLiA. This dataset collects PPG signals from 15 subjects for heart rate estimation while performing various daily activities. These activities include sitting, ascending/descending stairs, table soccer, cycling, driving, lunch break, walking, and working. As a result, the dataset captures a wide range of heart rates, varying from 60 to 150 beats per minute, depending on the specific activity being performed.
@@ D.1 Multi-class Classification
Table 9: Multi-class classification comparison against pre-trained models. The parameter size and reference are denoted next to their names. 95% CIs are reported in square brackets and the best value is bolded.
REGLE (0.07M) Chronos (200M) Moment (385M) PaPaGei-P (5M) PaPaGei-S (5.7M)
@@
Table 10: Multi-class classification comparison against CL methods. The parameter size and reference are denoted next to their names. 95% CIs are reported in square brackets and the best value is bolded.
Stat. Features SimCLR BYOL TF-C PaPaGei-P (5M) PaPaGei-S (5.7M)
@@
D.2 Other Evaluation Metrics
Table 11: Downstream comparison against pre-trained models (additional metrics: F1-score and $R^{2}$). The parameter size and reference are denoted next to their names. 95% CIs are reported in square brackets and the best value is bolded.
REGLE (0.07M) Chronos (200M) Moment (385M) PaPaGei-P (5M) PaPaGei-S (5.7M)
@@ D.2 Other Evaluation Metrics
Table 12: Downstream comparison against CL models (additional metrics: F1-score and $R^{2}$). The parameter size and reference are denoted next to their names. 95% CIs are reported in square brackets and the best value is bolded.
Stat. Features SimCLR BYOL TF-C PaPaGei-P (5M) PaPaGei-S (5.7M)

### v2

@@ 6 Discussion & Conclusion
Our results show that PaPaGei outperforms baselines in at least 14 tasks, with classification and regression improvements of 4.7%-6.3% and 2.9%-4.9%, respectively. PaPaGei-S excelled in cardiovascular tasks like BP, Hypertension, and HR, which can be attributed to the sVRI and IPA objectives, and PaPaGei-P outperformed baselines like Moment, excelling in tasks such as Smoking and Arousal. Ablation studies confirmed that the model with all three SSL objectives performs best, with sVRI highlighted as a key component and IPA and SQI providing positive knowledge transfer in multi-task setups. To assess performance under class imbalance, we examined the F1-score. PaPaGei achieves the highest F1 in 6 out of 9 classification tasks, demonstrating its effectiveness in handling data imbalance. For regression, PaPaGei-S achieves the highest $R^{2}$ in 7 tasks (Appendix §D), reflecting better alignment with the true distribution. These results highlight the robustness and versatility of PaPaGei-S across classification and regression tasks. PaPaGei is both data- and size-efficient (5M), making it ideal for medical applications where large models (200M+) are impractical due to on-device limitations or data privacy concerns with cloud model inference. While combining PaPaGei-P and PaPaGei-S objectives into one model might seem intuitive, it is impractical because it would constrain positive pairs on both sVRI and the number of participants, resulting in too many unique labels with limited samples per label. Our case studies also showed that PaPaGei-S captured personal medical information due to well-dispersed embeddings, compared to baselines. Future work should focus on diversifying training data, investigating sampling rate effects, and exploring multi-modal approaches or alternative architectures. Additionally, as extracting PPG features for different morphologies is non-trivial, future work benefit from systematic evaluation of PPG features and modeling. In conclusion, PaPaGei represents a significant advancement in foundation models for analyzing PPG signals in resource-constrained medical environments, with its open-source nature encouraging further research and development in healthcare applications.
@@
Models, data, and code are publicly available for reproducibility and future research.We exclusively utilize publicly accessible datasets, which can be requested or downloaded from the respective study group websites, allowing others to easily obtain the data for their own analyses. In §4 and Appendix §B, we provide comprehensive descriptions of the datasets, ground-truth annotations, and data pre-processing methods used in our experiments, ensuring transparency in our data handling procedures. The code to run our model is published with user-friendly examples. We have provided a detailed overview of the model architecture and its hyperparameters in §3, §4.1, and Appendix §A. Thus, our work is designed to be reproducible, enabling future research to build upon our findings.
@@ Appendix A Training and Inference Details
Architecture & Pre-training. The architecture of our ResNet 18-block encoder is described in Tables 7, 7, and 7. Each 1D convolution layer is configured with a kernel size of 3 and a stride of 2, while the max-pooling layer utilizes a kernel size of 3 with a stride of 1. We start with a filter size of 32, which doubles every 4 blocks to capture progressively more complex features. Dropout is applied with a probability of 0.5 to prevent overfitting. This backbone architecture is used across different methods in our experiments to ensure consistency and make for a fair comparison during evaluation. Additionally, our SSL baselines use the same batch size, learning rate, input sampling frequency, and training steps as PaPaGei. We use the same augmentation types and intensity for BYOL (Grill et al., 2020), SimCLR Chen et al. (2020), and PaPaGei-P. Furthermore, we investigated 0.07 and 0.5 temperatures as MoCo (He et al., 2020) and SimCLR (Chen et al., 2020), respectively. In contrast to a smaller embedding size of 256 adopted by (Abbaspourazad et al., 2023), we project the learned representations to a 512-dimensional embedding after the convolutional block (we investigated larger embedding sizes of 768 and 1024 and found no significant performance changes). This embedding is then passed through two Mixture of Experts (MoE) blocks, each containing three experts. Each expert block consists of two sequential linear layers, with sizes 256 and 1, which are used for IPA and SQI prediction tasks. It is noteworthy that both BYOL and TF-C require multiple encoders and different projection heads, resulting in variations in model sizes. For these methods, we use existing implementations available online [^3] https://github.com/chengding0713/SiamQuality [^4] https://github.com/mims-harvard/TFC-pretraining, but apply our encoder as the backbone to ensure consistency. Our models are pre-trained for 15,000 steps using the Adam optimizer, with a learning rate of $10^{-4}$. We use a batch size of 128 for training since after various trials we did not observe significant differences in performance with batch sizes of 64 and 256. We performed five iterations of pre-training and selected the best-performing model for each downstream task. For SimCLR and PaPaGei-P, a single model consistently achieves the best performance across all tasks. For BYOL, we select two models that perform best across all tasks. Similarly, for TF-C and PaPaGei-S, we choose three models with the highest performance. We use this approach as some models excel in certain task groups while others perform better in the rest. Note that a more robust approach would involve broader hyperparameter tuning with k-fold validation to obtain the optimal model. However, this requires substantial computational resources for pre-training. Additionally, we did not perform an exhaustive evaluation of different augmentation settings but instead used transformations and values based on prior research (Abbaspourazad et al., 2023; Tang et al., 2020). For model training, we primarily used PyTorch (Paszke et al., 2019). The NTXentLoss implementation was sourced from the PyTorch Metric Learning package [^5] https://github.com/KevinMusgrave/pytorch-metric-learning.
@@
Parameters: Training and Inference. This section outlines the training and inference parameters used in our methods. Inference parameters are those utilized for feature extraction.
  • PaPaGei-P (5M) and SimCLR (5M): Both training and inference involve 5M parameters. For SimCLR, it is worth noting that we use the projection features during inference instead of using the encoder only.
  • BYOL (5M): During training, the online and target encoders each have 5M parameters, and the projector is 800K. At inference, only the online encoder is used for feature extraction, totaling 5M parameters.
  • TF-C (10M): The time and frequency encoders each have 5M parameters, followed by a smaller projector ($<$ 100K). Since both encoders and projectors are required for inference, the total parameter count is 10M.
  • PaPaGei-S (5M): The encoder consists of 5M parameters, while the expert heads contribute approximately 400K each. As the expert heads are not used for feature extraction, the inference parameter total remains 5M.
@@ Appendix B Datasets and Tasks
Table 8: The task evaluation benchmark of PaPaGei. Datasets highlighted in gray are unseen during training, thus, the corresponding tasks are out-of-domain. The rest were used for pre-training but their test sets and labels are held out. For task type, B/M/R refer to Binary classification, Multi-class classification (#classes), and Regression, respectively.
#ID Dataset SR (Hz) Collected by Task Task Type #Participants (#Samples)
@@
VitalDB. The VitalDB dataset provides comprehensive monitoring of vital signs and physiological parameters from 6,388 surgical cases. This high-resolution dataset includes a wide range of intraoperative monitoring variables such as heart rate, blood pressure, oxygen saturation, and other critical physiological signals, collected at frequent intervals throughout surgery. The surgical operation belongs to one of the eleven categories: colorectal, biliary/pancreas, stomach, major resection, minor resection, breast, transplantation, thyroid, hepatic, vascular, and others. After the data cleaning process, we narrowed the dataset down to 5,866 participants with complete and usable information. As depicted in Figure 10, we observe that the gender distribution is relatively balanced, with nearly equal representation of male and female patients. Additionally, the majority of the participants fall within the age range of 50 to 70, with a significant proportion being around 60 years old. The ICU label corresponds to whether the person was admitted to the ICU or not.
@@
MIMIC-III. In our analysis, we utilize the MIMIC-III waveform database matched subset, which comprises data from 10,282 ICU patients. From this dataset, we focus specifically on extracting photoplethysmogram (PPG) data, provided it is available for each patient. To ensure the quality of the data, we set a criterion of at least 1 minute of usable PPG signal that must be present. After performing a thorough data cleaning process, we end up with a cohort of 5,596 participants with reliable PPG data. As illustrated in Figure 11, the dataset shows a gender imbalance, with a higher proportion of male patients compared to female patients. Additionally, the majority of participants are aged 60 years or older, reflecting a typical ICU population that often includes elderly patients with critical health conditions.
@@
MESA. The Multi-Ethnic Study of Atherosclerosis (MESA) sleep sub-study gathered data from 2,237 participants through overnight, unattended polysomnography to assess various sleep parameters. After the data cleaning process, we retained 2,055 participants for analysis. As shown in Figure 12, the dataset shows a slightly larger proportion of female participants. The age distribution reveals that most participants are between 60 and 80 years old, reflecting an older adult population, which is commonly studied concerning sleep disorders and cardiovascular risks.
In this study, we use the Apnea-Hypopnea Index (AHI) with at least 3% and 4% oxygen desaturation as the primary measure for diagnosing sleep apnea, as recommended by the American Academy of Sleep Medicine (Ruehland et al., 2009). These thresholds indicate the severity of sleep apnea, with oxygen desaturation during apneas/hypopneas being a critical factor. We predict these AHI values directly in our regression models. Additionally, we classify participants with any history of smoking as smokers. This approach allows us to account for both current and former smokers, capturing a broader range of smoking-related health risks within our analysis.
@@
VitalVideos (VV) (Skin Tone). The Vital Videos study is an ongoing project that collects data on vital signs, videos, and blood pressure across a variety of conditions, including variations in lighting, background, and skin tone. For our analysis, we used data from two groups, totalling 231 participants, from Europe and Sub-Saharan Africa. As shown in Figure 14, most participants have a Fitzpatrick skin tone of 5 or 6, indicating darker skin. The dataset is primarily composed of female participants, with an age range between 40 and 60 years. Additionally, the majority of participants had a systolic blood pressure of around 125 and a diastolic pressure of around 80, suggesting that most individuals in the study were relatively healthy.
@@
PPG-BP. The PPG-BP consists of short PPG recordings from 219 participants collected at 1000Hz. For each subject, there are three 2.1s recordings. For our analysis, we zero pad them to 10s. In Figure 15, the age distribution shows that most participants are between 40 and 80 years old, with fewer participants under 40. Furthermore, the majority of individuals have hypertension. In terms of gender, the dataset has slightly more females than males. The distribution of systolic blood pressure is centered around 120-140, indicating a population with normal to moderately elevated blood pressure, while diastolic blood pressure predominantly falls between 70 and 90. Lastly, the average heart rate for most participants ranges between 70 and 90 beats per minute.
@@
ECSMP. The ECSMP dataset was gathered to study the relationship between emotion, cognition, and sleep in 89 participants. As shown in Figure 17, the majority of the participants are young adult females, with an average age of around 25 years. Mood disturbances were measured using the Profile of Mood States (POMS) scale, which captures various aspects of emotional states. To classify participants into high versus low mood disturbance categories, we binarized the Total Mood Disturbance (TMD) values by using the median as the cutoff point.
@@
WESAD. The wearable stress and affect detection dataset is a multi-modal dataset collected from 15 participants using various sensor modalities. In this study, participants were exposed to videos designed to elicit different affective states, such as amusement, meditation, stress, and baseline conditions. Following each session, participants completed the Self-Assessment Manikins (SAM) questionnaire (Bradley & Lang, 1994), which provided the ground-truth values for valence and arousal. In our analysis, we binarized these values by categorizing valence and arousal as low (1) when less than 5 and high (0) otherwise. Then, we perform regression at the segment level. As shown in Figure 18, arousal levels are generally low, while valence tends to be high in most cases.
@@
PPG-DaLiA. This dataset collects PPG signals from 15 participants for heart rate estimation while performing various daily activities. These activities include sitting, ascending/descending stairs, table soccer, cycling, driving, lunch break, walking, and working. As a result, the dataset captures a wide range of heart rates, varying from 60 to 150 beats per minute, depending on the specific activity being performed. To align the PPG signal with the activity labels, we use a 8s window with 6s and 2s overlap and shift, respectively. After this, we resample and pad the signal to facilitate modeling.
@@ D.1 Multi-class Classification
Table 9: Multi-class classification comparison against pre-trained models. Feature extraction parameters are indicated next to each name.. 95% CIs are reported in square brackets and the best value is bolded.
REGLE (0.07M) Chronos (200M) Moment (385M) PaPaGei-P (5M) PaPaGei-S (5M)
@@
Table 10: Multi-class classification comparison against CL methods. Feature extraction parameters are indicated next to each name.. 95% CIs are reported in square brackets and the best value is bolded.
Stat. Features SimCLR (5M) BYOL (5M) TF-C (10M) PaPaGei-P (5M) PaPaGei-S (5M)
@@
D.2 F1-Score and $R^{2}$ evaluation metrics.
Table 11: Downstream comparison against pre-trained models (additional metrics: F1-score and $R^{2}$). Feature extraction parameters are indicated next to each name. 95% CIs are reported in square brackets and the best value is bolded.
REGLE (0.07M) Chronos (200M) Moment (385M) PaPaGei-P (5M) PaPaGei-S (5M)
@@
Table 12: Downstream comparison against CL models (additional metrics: F1-score and $R^{2}$). Feature extraction parameters are indicated next to each name. 95% CIs are reported in square brackets and the best value is bolded.
Stat. Features SimCLR (5M) BYOL (5M) TF-C (10M) PaPaGei-P (5M) PaPaGei-S (5M)
@@
D.3 Ablation Results
In this section, we provide the numeric results for the scaling analysis (Table 13) and PaPaGei-S component analysis (Table 14).